OpenStudy (anonymous):

The number of people in a small city of 4000 residents who became ill with a particular virus, t weeks its initial outbreak, is described by the function f(t)=40000/1+30e^1.2t How many people were ill at the end of the fourth week?

OpenStudy (anonymous):

just put t=4 in the function \[f(t)=40000/1+30e^1.2(4)\]

OpenStudy (anonymous):

f(t)=40000/1+30e^ 1.2(4) f(t)=40000 +3645.31 f(t)=43645.31

OpenStudy (anonymous):

im very confused

OpenStudy (anonymous):

what u don't understand ?!

OpenStudy (anonymous):

fixed explanation 1.you have the function call f(t)=40000/1+30e^1.2t 2.f(t) means f(at t which is number of week) 3.to get the fourth week u need to put 4 inside the function like this f(4)=40000/1+30e^1.2(4) 4.calculate the numbers f(4)=40000/1+30e^1.2(4) f(4)=40000/1+30e^(4.8) f(4)=40000/1+30*(121.51) f(4)=40000/1+3645.3 f(4)=40000/3646.3 f(4)=10.9700 since it is calculate people u cant get Fraction so answer is f(4)=11

OpenStudy (anonymous):

\[\large f(t)=\frac{40000}{1+30e^{1.2t}}\] \[\large f(4)=\frac{40000}{1+30e^{1.2\times 4}}\] and a calculator